OpenStudy (anonymous):

Which answer shows this equation written in standard form? 9 – 4(x – y) = –6x + 6 A. –10x + 4y = –3 B. 2x + 4y = 3 C. 2x – 4y = 3 D. 2x + 4y = –3

OpenStudy (anonymous):

im really stupid i need help

OpenStudy (anonymous):

lol

OpenStudy (michele_laino):

we have: left side= \[9-4x+4y\] then I rewrite your eqaution as below: \[9-4x+4y=-6x+6\]

OpenStudy (anonymous):

so know do i add the 4's and 6's together

OpenStudy (michele_laino):

now I add to both sides 6x, namely: \[9-4x+4y+6x=-6x+6+6x\] so: \[2x+4y-9=6\] Now I subtract to both sides 9, namely: \[2x+4y+9-9=6-9\] so: \[2x+4y=-3\] so D.

OpenStudy (anonymous):

thank you so much im so stupid

OpenStudy (michele_laino):

I think you are intelligent, you only have to study much more! @zlamar11

OpenStudy (michele_laino):

thank you! @zlamar11